# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation On Diverse Modalities

### v3

@@ E.4 F1-score of benchmark
Mapping CORAL 0.74 0.76 0.74 0.16 0.65 0.57 0.62 0.72 0.65 0.72 0.92 0.62 CircV 5.00
@@ E.6 Comparisons between supervised and unsupervised scorers
We observe in Figure 6 the cross-validation score as a function of the final accuracy for various DA methods type and for both supervised and unsupervised scorers. As expected, we observe a good correlation between accuracy and cross-validation score with the supervised scorer. An important remark is that the Circular Validation (CircV) (Bruzzone & Marconcini, 2010b) shows some correlation between accuracy and cross-validation score. It indicates that this unsupervised scorer might be the most suitable choice for hyperparameter selection. This is supported by our extended experimental results in Table 2 for which the CircV is selected as the best scorer the most often. A similar trend can be observed for the Importance Weighted (IW) (Sugiyama et al., 2007a) which is also confirmed in Table 2.

### v4

@@ E.4 F1-score of benchmark
Mapping
CORAL 0.74 0.76 0.74 0.16 0.65 0.57 0.62 0.72 0.65 0.72 0.92 0.62 CircV 5.00
@@ E.6 Comparisons between supervised and unsupervised scorers
We observe in Figure 7 the cross-validation score as a function of the final accuracy for various DA methods type and for both supervised and unsupervised scorers. As expected, we observe a good correlation between accuracy and cross-validation score with the supervised scorer. An important remark is that the Circular Validation (CircV) (Bruzzone & Marconcini, 2010b) shows some correlation between accuracy and cross-validation score. It indicates that this unsupervised scorer might be the most suitable choice for hyperparameter selection. Using spearman correlation shows the same conclusion (see Figure 6). This is supported by our extended experimental results in Table 2 for which the CircV is selected as the best scorer the most often. A similar trend can be observed for the Importance Weighted (IW) (Sugiyama et al., 2007a) which is also confirmed in Table 2.
Figure 6: Cross-val score as a function of the accuracy for different supervised and unsupervised scorers. The Spearman correlation coefficient is reported for each scorer by $\rho$. Each point represents an inner split with a DA method (color of the points) and a dataset. A good score should correlate with the target accuracy.